OpenStudy (anonymous):

solve the following system of equations. 2x=4 2x-y=5 i will give a medal

OpenStudy (anonymous):

2x = 4 x = 2 plug in x in 2x - y = 5 and solve for y

OpenStudy (anonymous):

2x - y = 5 2(2) - y = 5 4 - y = 5 -y = 1 y = -1 (2,-1) Im sorry that isnt one of your answer choices, did you type the equations right in your question?

OpenStudy (anonymous):

Yes can u help me with 2 more plz

OpenStudy (anonymous):

sure, post em

OpenStudy (anonymous):

find the y-intercept of 5x+2y=10 My answers include: (0,2), (0,5), (2,0), (5,0) is the first one

OpenStudy (anonymous):

the 2nd one is: which of the points is not a solution to -4x-2y=8? My answers include: (-2,0), (0,-4), (1,-6), (1,2)

OpenStudy (anonymous):

number 1) set x to 0 and solve for y number 2) plug in the values for x and y. if the answers do not match (such as 8 = 8) then it is not inclided

OpenStudy (anonymous):

confused??

OpenStudy (anonymous):

number 1: 5x+2y=10 5(0) + 2y = 10 solve for y in the above equation number 2: you know the coordinates are (x,y) right? I will use the coordinates (-2,0) as an example -4x - 2y = 8 -4(-2) - 2(0) = 8 solve and repeat with other coordinates

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

no problem